# Robust estimation of boundary using doubly stochastic scaling of Gaussian kernel

### v2

@@ 2. Characterizing doubly stochastic scaling factors on manifolds with boundary
(24) $\rho_{h}(\mathbf{x})^{2}q(\mathbf{x})=1+\mathcal{O}(h^{2}).$
Moreover, the constant in $\mathcal{O}(h^{2})$ that depends on $\mathbf{x}$ can be uniformly bounded since $\mathcal{M}$ is compact.
In fact, for a sufficiently small $h$ the above result holds within the interior of a manifold with nonempty boundary where $b_{\mathbf{x}}/h\gg 0$. The key component of the proof is the asymptotic expansion of the heat kernel in the interior of the manifold [25, 9], combined with the fact that when $b_{\mathbf{x}}/h\rightarrow\infty$ then $m^{(0)}_{h}(\mathbf{x})\rightarrow\pi^{d/2}$, $m^{(1)}_{h}(\mathbf{x})\rightarrow 0$, $\frac{b_{\mathbf{x}}}{h}m^{(1)}_{h}(\mathbf{x})\rightarrow 0$ and $m^{(2)}_{h}(\mathbf{x})\rightarrow\pi^{d/2}/2$. Consequently, $A_{h}(\mathbf{x})\rightarrow-\pi^{d}$ and $B_{h}(\mathbf{x})\rightarrow 0$ in Eq. 21. Thus $\rho_{h}(\mathbf{x})^{4}q(\mathbf{x})^{2}\rightarrow 1+\mathcal{O}(h^{2})$.
In addition to the cases of closed manifolds and the interior of a manifold with a boundary, a scenario where a second-order approximation of $\rho_{h}(\mathbf{x})^{2}q(\mathbf{x})$ can be easily described arises when $H(\mathbf{x})=0$, as with a flat boundary, and $\partial_{\bm{\eta}_{\mathbf{x}}}q(\mathbf{x})=0$, as with a uniform sampling density. This is detailed in the corollary below.
@@
If $\partial\mathcal{M}\neq\emptyset$ then for $\mathbf{x}\in\partial\mathcal{M}$ (i.e. $b_{\mathbf{x}}=0$),
(26) $A_{h}(\mathbf{x})$ $=\frac{\pi^{d-1}}{2}\left(1-\frac{\pi}{2}\right)+\frac{h(d-1)\pi^{d-1/2}H(\mathbf{x})}{8}$
(27) $B_{h}(\mathbf{x})$ $=-\frac{\pi^{d-1/2}}{2}h\omega(\mathbf{x}).$
Therefore, for practical purposes, we relax the result in Theorem 2.1 to obtain a first-order characterization of $\rho_{h}(\mathbf{x})^{2}q(\mathbf{x})$ that depends exclusively on $b_{\mathbf{x}}$, removing the reliance on $H(\mathbf{x})$ and $\partial_{\bm{\eta}_{\mathbf{x}}}q(\mathbf{x})$, that are not accessible for a given dataset.
@@
is a strictly decreasing convex function with a range of $\left[2\sqrt{\frac{\pi}{\pi-2}},1\right)$.
The above result will be utilized in the convergence proof of our doubly stochastic kernel-based BDE presented in Theorem 3.1 in the following section.
3. Estimation of boundary using doubly stochastic kernel
In this section, we first review the standard Gaussian kernel-based BDE proposed in [1] and how it is utilized to estimate the boundary points. Next, we propose a BDE that leverages the doubly stochastic kernel obtained through Sinkhorn iterations [17, 10] applied to the Gaussian kernel, and local PCA. Using Theorem 2.1 and Corollary 2.4 we establish the convergence of our BDE to the normal direction to the boundary, up to a factor that depends only on the distance to the boundary. Building on this result, we introduce a straightforward algorithm for computing boundary points of the given data manifold.
@@ 3.1. Boundary estimation using standard Gaussian kernel
Using the standard Gaussian kernel $\mathbf{K}_{ij}$ as defined in Eq. 6, Berry and Sauer [1] defined a KDE $f$ and a BDE $\bm{\mu}$ as follows,
@@
Intuitively, $\bm{\mu}_{i}$ vanishes in the interior while, near the boundary, it is non-zero and points in the direction of $-\bm{\eta}_{\mathbf{x}_{i}}$, as illustrated in Figure 1.
@@
Berry and Sauer [1] showed in the noiseless setting that $\bm{\mu}_{i}$ converges to the normal direction $-\bm{\eta}_{\mathbf{x}_{i}}$ up to a global factor that depends on the distance $b_{\mathbf{x}_{i}}$ of $\mathbf{x}_{i}$ from the boundary and the sampling density $q(\mathbf{x}_{i})$ at $\mathbf{x}_{i}$. They also showed that $f_{i}$ converges to the sampling density $q(\mathbf{x}_{i})$ up to another global factor that depends on $b_{\mathbf{x}_{i}}$. Precisely,
@@
They finally estimated the distances to the boundary $\widehat{b}_{\mathbf{x}_{i}}$ by finding the root of the equation $F(b_{\mathbf{x}})=hf_{i}/\sqrt{\pi}\left\|\mathbf{\mu}_{i}\right\|$ using Newton’s method where $F(b_{\mathbf{x}})=\left(1+\mathrm{erf}\left(\frac{b_{\mathbf{x}}}{h}\right)\right)e^{b_{\mathbf{x}}^{2}/h^{2}}$.
@@ 3.2. Boundary estimation using doubly stochastic kernel and local PCA
We propose a BDE by combining the the doubly stochastic kernel [17] and local PCA as follows,
@@
Here, $\widetilde{\mathbf{U}}_{i}\in\mathbb{R}^{m\times d}$ serves as an approximation of an orthogonal basis of $T_{\mathbf{x}_{i}}\mathcal{M}$ and is obtained by applying PCA on a neighborhood of $\widetilde{\mathbf{x}}_{i}$. Unlike the Gaussian kernel-based approach [1], which utilizes vectors $\widetilde{\mathbf{x}}_{j}-\widetilde{\mathbf{x}}_{i}$ between the original (potentially noisy) data points, our method employs vectors between the PCA-projected points, $\widetilde{\mathbf{U}}_{i}^{T}(\widetilde{\mathbf{x}}_{j}-\widetilde{\mathbf{x}}_{i})$. At a high level, this design leverages the noise-resilient properties of the doubly stochastic kernel [17, 8] combined with the denoising capabilities of local PCA [11].
@@
    The projector onto the tangent space is assumed to be available. Specifically, we assume access to the matrices $\{\bm{\mathcal{U}}_{i}\}_{1}^{n}\subset\mathbb{R}^{m\times d}$, where the columns of $\bm{\mathcal{U}}_{i}$ form an orthonormal basis of $T_{\mathbf{x}_{i}}\mathcal{M}$. In our experiments, however, we approximate an orthonormal basis of $T_{\mathbf{x}_{i}}\mathcal{M}$ using $\widetilde{\mathbf{U}}_{i}$, obtained by applying local PCA to noisy observations. For readers interested in the analysis showing that $\widetilde{\mathbf{U}}_{i}$ is close to a true basis $\bm{\mathcal{U}}_{i}\mathbf{S}_{i}$ for some $\mathbf{S}_{i}\in\mathbb{O}(d)$, we refer to the following papers [22, 24, 7, 27].
@@
    The noise $\bm{\varepsilon}(\mathbf{x}_{i})$ is orthogonal to the tangent space i.e. $\bm{\mathcal{U}}_{i}^{T}\bm{\varepsilon}(\mathbf{x}_{i})=0$.
With assumptions (A6) and (A7), we are now ready to state our main result, which characterizes our BDE $\bm{\nu}_{i}$.
@@
(36) $\bm{\nu}_{i}=h\beta_{h}(\mathbf{x})\bm{\eta}_{\mathbf{x}}+\mathcal{O}\left(h^{2}\right)+\mathcal{O}^{(h)}(g(m,n))$
where $\mathcal{O}^{(h)}(g(m,n))$ is defined in Theorem 1.1 and
(37) $\beta_{h}(\mathbf{x})=\frac{m_{h}^{(1)}(\mathbf{x})\zeta_{h}(\mathbf{x})}{\pi^{d/2}+2m_{h}^{(2)}(\mathbf{x})\zeta_{h}(\mathbf{x})}.$
Here, $-\beta_{h}(\mathbf{x})$ is a positive and strictly decreasing convex function that depends only on $b_{\mathbf{x}}$, the distance from $\mathbf{x}$ from the boundary, and decays from $\frac{1}{2}(\sqrt{\pi}-\sqrt{\pi-2})$ to 00 as $b_{\mathbf{x}}$ goes from 00 to $\infty$.
@@
Under assumptions (A1, A2’, A3-A7), if $q$ is uniform i.e. $q=1/|\mathcal{M}|$, we have
(38) $\bm{\nu}_{i}=h\beta_{h}(\mathbf{x})\bm{\eta}_{\mathbf{x}}+\mathcal{O}\left(h^{3}\right)+\mathcal{O}^{(h)}(g(m,n)).$
The proof follows directly from the proof of Theorem 3.1 by substituting zeros for the partial derivatives of $q$. From the above results, it follows that
@@
for sufficiently small $h$, with high probability in the large sample and high-dimensional setting. Since $\beta_{h}(\mathbf{x})$ depends solely on $b_{\mathbf{x}}$ and is strictly decreasing, we can identify the boundary points by simply thresholding $\{\left\|\bm{\nu}_{i}\right\|\}_{1}^{n}$, for example by using a chosen percentile of these values. The steps to compute $(\bm{\nu}_{i})_{1}^{n}$ are summarized in Algorithm 1, and it differs from the Gaussian kernel-based approach described in Section 3.1 in two ways: (i) it employs Sinkhorn iterations [10, 17] to construct the doubly stochastic kernel, and (ii) it eliminates the need for KDE computations and Newton’s method to approximate boundary points.
Algorithm 1 Boundary direction estimator
0:  $(\widetilde{\mathbf{x}}_{i})_{1}^{n}\subset\mathbb{R}^{m}$, $d$ $h$, $k_{\text{nn}}$.
@@
2:  Construct $\mathbf{W}$ as in Eq. (7) using Sinkhorn iterations [10, 17].

### v3

@@ 3. Characterizing doubly stochastic scaling factors on manifolds with boundary
$\rho_{h}(\mathbf{x})^{2}q(\mathbf{x})=1+\mathcal{O}(h^{2})$
where the constant in $\mathcal{O}(h^{2})$ which depends on $\mathbf{x}$ can be uniformly bounded since $\mathcal{M}$ is compact.
In fact, for a sufficiently small $h$ the above result holds at the interior points of a manifold with nonempty boundary whenever $b_{\mathbf{x}}/h\gg 0$. The key component of the proof is the asymptotic expansion of the heat kernel in the interior of the manifold [vaughn2019diffusion, coifman2006diffusion], combined with the fact that when $b_{\mathbf{x}}/h\rightarrow\infty$ then $m^{(0)}_{h}(\mathbf{x})\rightarrow\pi^{d/2}$, $\frac{b_{\mathbf{x}}}{h}m^{(1)}_{h}(\mathbf{x})\rightarrow 0$ and $m^{(2)}_{h}(\mathbf{x})\rightarrow\pi^{d/2}/2$. Consequently, $A_{h}(\mathbf{x})\rightarrow-\pi^{d}$ and $B_{h}(\mathbf{x})\rightarrow 0$ in Eq. 24, which yields $\rho_{h}(\mathbf{x})^{4}q(\mathbf{x})^{2}\rightarrow 1+\mathcal{O}(h^{2})$.
In addition to the cases of closed manifolds and the interior of a manifold with a boundary, a scenario where a second-order approximation of $\rho_{h}(\mathbf{x})^{2}q(\mathbf{x})$ can be easily described arises when $H(\mathbf{x})=0$, as with a flat boundary, and $\partial_{\bm{\eta}_{\mathbf{x}}}q(\mathbf{x})=0$, as with a uniform sampling density.
@@
If $\partial\mathcal{M}\neq\emptyset$ then for $\mathbf{x}\in\partial\mathcal{M}$,
(28) $A_{h}(\mathbf{x})$ $=\frac{\pi^{d-1}}{2}\left(1-\frac{\pi}{2}\right)+\frac{h(d-1)\pi^{d-1/2}H(\mathbf{x})}{4}$
(29) $B_{h}(\mathbf{x})$ $=-\frac{\pi^{d-1/2}}{2}h\kappa_{+}(\mathbf{x}).$
Therefore, for practical purposes, we relax the result in Theorem 3.1 to obtain a first-order characterization of $\rho_{h}(\mathbf{x})^{2}q(\mathbf{x})$ that depends exclusively on $b_{\mathbf{x}}$, removing the reliance on $H(\mathbf{x})$ and $\partial_{\bm{\eta}_{\mathbf{x}}}q(\mathbf{x})$, which are typically not accessible for a given dataset.
@@
is a strictly decreasing convex function that depends only on $b_{\mathbf{x}}$, the distance of $\mathbf{x}$ from the boundary, with a range of $\left(1,2\sqrt{\frac{\pi}{\pi-2}}\right]$.
The above characterization of the scaling factors $\rho_{h}(\mathbf{x})$ is utilized in the convergence analysis of our doubly stochastic kernel-based BDE, which is presented in the next section.
4. Estimation of boundary and sampling density using doubly stochastic kernel
In this section, we first review the standard Gaussian kernel-based BDE proposed in [berry2017density] and how it is utilized to estimate the boundary points. Next, we propose a BDE that leverages the doubly stochastic kernel obtained via Sinkhorn iterations [landa2023robust, cuturi2013sinkhorn] applied to the Gaussian kernel, and local PCA. Using Theorem 3.1 and Corollary 3.4 we establish the convergence of our BDE to the normal direction to the boundary, up to a factor that depends only on the distance to the boundary. Building on this result, we introduce a straightforward algorithm for computing boundary points of the given data manifold. Finally, we propose a two-parameter family of KDEs that utilize the doubly stochastic kernel and converge to the sampling density up to a scaling factor that depends on the distances to the boundary.
@@ 4.1. Boundary estimation using standard Gaussian kernel
Using the standard Gaussian kernel $\mathbf{K}_{ij}$ as defined in Eq. 7, Berry and Sauer [berry2017density] defined a KDE $f$ and a BDE $\bm{\mu}$ as follows,
@@
Intuitively, $\bm{\mu}_{i}$ vanishes in the interior while, near the boundary, it is non-zero and points in the direction of $\bm{\eta}_{\mathbf{x}_{i}}$, i.e. opposite to the normal direction to the boundary, as illustrated in Figure 1.
@@
Berry and Sauer [berry2017density] showed in the noiseless setting that $\bm{\mu}_{i}$ converges to the normal direction $\bm{\eta}_{\mathbf{x}_{i}}$ up to a global factor that depends on the distance $b_{\mathbf{x}_{i}}$ of $\mathbf{x}_{i}$ from the boundary and the sampling density $q(\mathbf{x}_{i})$ at $\mathbf{x}_{i}$. They also showed that $f_{i}$ converges to the sampling density $q(\mathbf{x}_{i})$ up to another global factor that depends on $b_{\mathbf{x}_{i}}$. Precisely,
@@
The distances to the boundary $\widehat{b}_{\mathbf{x}_{i}}$ are then estimated by finding the root of the equation $F(b_{\mathbf{x}})=hf_{i}/\sqrt{\pi}\left\|\bm{\mu}_{i}\right\|$ using Newton’s method where $F(b_{\mathbf{x}})=\left(1+\mathrm{erf}\left(\frac{b_{\mathbf{x}}}{h}\right)\right)e^{b_{\mathbf{x}}^{2}/h^{2}}$.
@@ 4.2. Boundary estimation using doubly stochastic kernel and local PCA
Here, we propose a BDE by replacing the standard Gaussian kernel $\mathbf{K}$ in Eq. (33) with the doubly stochastic kernel $\mathbf{W}$ as defined in Eq. (8). Moreover, instead of using the vectors $\widetilde{\mathbf{x}}_{j}-\widetilde{\mathbf{x}}_{i}$ between the original (potentially noisy) data points, our method employs vectors between the PCA-projected points, $\widetilde{\mathbf{U}}_{i}^{T}(\widetilde{\mathbf{x}}_{j}-\widetilde{\mathbf{x}}_{i})$ where $\widetilde{\mathbf{U}}_{i}\in\mathbb{R}^{m\times d}$ serves as an approximation of an orthogonal basis of $T_{\mathbf{x}_{i}}\mathcal{M}$ and is obtained by applying PCA on a neighborhood of $\widetilde{\mathbf{x}}_{i}$. At a high level, this design leverages the noise-resilient properties of the doubly stochastic kernel [landa2023robust, cheng2024bi] combined with the denoising capabilities of local PCA [gong2010locally]. Overall, our BDE at the $i$-th point is given by,
@@
    The projector onto the tangent space is assumed to be available. Specifically, we assume access to the matrices $\{\bm{\mathcal{U}}_{i}\}_{1}^{n}\subset\mathbb{R}^{m\times d}$, where the columns of $\bm{\mathcal{U}}_{i}$ form an orthonormal basis of $T_{\mathbf{x}_{i}}\mathcal{M}$. In our experiments, however, we approximate an orthonormal basis of $T_{\mathbf{x}_{i}}\mathcal{M}$ using $\widetilde{\mathbf{U}}_{i}$, obtained by applying local PCA to noisy observations. For readers interested in the analysis showing the convergence of $\widetilde{\mathbf{U}}_{i}$ to a true basis $\bm{\mathcal{U}}_{i}\mathbf{S}_{i}$ for some $\mathbf{S}_{i}\in\mathbb{O}(d)$, we refer to the following papers [singer2012vector, tyagi2013tangent, cheng2013local, kaslovsky2014non, yu2015useful, mohammed2017manifold, aamari2019nonasymptotic].
@@
    The noise $\bm{\varepsilon}(\mathbf{x}_{i})$ is orthogonal to the tangent space, i.e. $\bm{\mathcal{U}}_{i}^{T}\bm{\varepsilon}(\mathbf{x}_{i})=0$.
With assumptions (A6) and (A7), we are now ready to state our main result, which characterizes our BDE $\bm{\nu}_{i}$. The following theorem establishes that our BDE using a doubly stochastic kernel and local PCA converges to the true normal direction $\bm{\eta}_{\mathbf{x}_{i}}$ up to a scaling factor that depends only on the distance from the boundary.
@@
(38) $\bm{\nu}_{i}=h\beta_{h}(\mathbf{x}_{i})\bm{\eta}_{\mathbf{x}_{i}}+\mathcal{O}\left(h^{2}\right)+\mathcal{O}^{(h)}(g(m,n))$
where $\mathcal{O}^{(h)}(g(m,n))$ is defined in Lemma 2.1 and
(39) $\beta_{h}(\mathbf{x}_{i})=\frac{m_{h}^{(1)}(\mathbf{x}_{i})\zeta_{h}(\mathbf{x}_{i})}{\pi^{d/2}+2m_{h}^{(2)}(\mathbf{x}_{i})\zeta_{h}(\mathbf{x}_{i})}.$
Here, $\beta_{h}(\mathbf{x})$ is a positive and strictly decreasing convex function that depends only on $b_{\mathbf{x}}$, the distance of $\mathbf{x}$ from the boundary, and decays from $\frac{1}{2}(\sqrt{\pi}-\sqrt{\pi-2})$ to 0 as $b_{\mathbf{x}}\rightarrow\infty$.
@@
Under assumptions (A1, A2’, A3-A7), if $H(\mathbf{x}_{i})=0$ and $q$ is uniform, then
$\bm{\nu}_{i}=h\beta_{h}(\mathbf{x}_{i})\bm{\eta}_{\mathbf{x}_{i}}+\mathcal{O}\left(h^{3}\right)+\mathcal{O}^{(h)}(g(m,n)).$
The proof follows directly from the proof of Theorem 4.1 by substituting zeros for $H$ and the partial derivatives of $q$.
From the above results, it follows that
@@
for sufficiently small $h$, with high probability in the large sample and high-dimensional setting. Since $\beta_{h}(\mathbf{x})$ depends solely on $b_{\mathbf{x}}$ and is strictly decreasing, we can identify the boundary points by simply thresholding $\{\left\|\bm{\nu}_{i}\right\|\}_{1}^{n}$, for example by using a chosen percentile of these values. The steps to compute $(\bm{\nu}_{i})_{1}^{n}$ are summarized in Algorithm 1, and it differs from the Gaussian kernel-based approach [berry2017density] described in Section 4.1 in two ways: (i) it employs Sinkhorn iterations [cuturi2013sinkhorn, landa2023robust] to construct the doubly stochastic kernel, and (ii) it eliminates the need for KDE computations and Newton’s method to approximate boundary points. In summary, normalizing the kernel via doubly stochastic scaling reduces the impact of noise [landa2021doubly, cheng2024bi], while local PCA provides additional noise filtering, improving the robustness of the estimates. Our method also maintains computational efficiency by identifying boundary points through a straightforward thresholding of the estimated boundary direction’s norm.
Algorithm 1 Boundary direction and points estimator
[…]
@@
2: Construct $\mathbf{W}$ as in Eq. (8) using Sinkhorn iterations [cuturi2013sinkhorn, landa2023robust].
@@
[…]
Having detected the boundary points, we can easily estimate distances to the boundary by computing the minimum of the shortest path distances of each point from the boundary points. To ensure robustness to noise, instead of using naive Euclidean distances between nearest neighbors as input to the shortest path algorithm, we use Euclidean distances computed from the PCA-projected embeddings of local neighborhoods. The pseudocode is provided below.
Algorithm 2 Distance to boundary estimator
[…]
1:for $i=1\rightarrow n$ do
2:  $\widetilde{\mathcal{N}}_{i}\leftarrow k_{\text{nn}}$-nearest neighbors of $\widetilde{\mathbf{x}}_{i}$.
3:  $\widetilde{\mathbf{U}}_{i}\leftarrow$ PCA on $\{\widetilde{\mathbf{x}}_{j}:j\in\widetilde{\mathcal{N}}_{i}\}$.
4:  $\mathrm{dist}_{ij}\leftarrow\left\|\widetilde{\mathbf{U}}_{i}^{T}(\widetilde{\mathbf{x}}_{j}-\widetilde{\mathbf{x}}_{i})\right\|_{2}$ for all $j\in\widetilde{\mathcal{N}}_{i}$.
5:end for
[…]
4.3. Density estimation using doubly stochastic kernel
Standard kernel density estimators for distributions with bounded support are well known to exhibit bias near the boundary [jones1996simple, marron1994transformations, cowling1996pseudodata, karunamuni2005boundary]. In particular, without explicit correction, kernel mass extends beyond the boundary of the support, leading to systematic underestimation of the density in regions close to the boundary. This phenomenon also persists in the manifold setting. However, having identified the boundary in the previous section, we can utilize the distance to the boundary to construct an explicit correction. Here, we introduce a two-parameter family of kernel density estimators that extends the method of [landa2023robust], originally developed for closed manifolds, by incorporating a boundary correction term based on $m^{(0)}_{h/\sqrt{s}}(\mathbf{x})$ and $\zeta_{h}(\mathbf{x})$, as defined in Eq. (19, 31). The resulting estimator is given by:
(41) $\widehat{q}_{i}\coloneqq\frac{1}{(n-1)h^{d}}\left(\frac{1}{\pi^{ds/2}s^{d/2}}\frac{1}{(m^{(0)}_{h/\sqrt{s}}(\mathbf{x}_{i})\zeta_{h}(\mathbf{x}_{i})^{s})^{u}}\sum_{j=1}^{n}\frac{\mathbf{W}_{ij}^{s}}{(m^{(0)}_{h/\sqrt{s}}(\mathbf{x}_{j})\zeta_{h}(\mathbf{x}_{j})^{s})^{1-u}}\right)^{1/(1-s)}$
where $s>0,s\neq 1$ and $u\in(0,1)$. In the absence of the terms involving $m^{(0)}_{h/\sqrt{s}}$ and $\zeta_{h}$, the above formulation reduces to the estimator in [landa2023robust] for closed manifolds (up to a global constant that depends on $h$, $s$ and $d$). The following theorem establishes the convergence of the above KDE to the true sampling density.
Theorem 4.2.
Under assumptions (A1, A2’, A3-A5), we have
(42) $\widehat{q}_{i}=\left(q(\mathbf{x}_{i})+\mathcal{O}(h)\right)(1+\mathcal{O}^{(h)}(g(m,n))),$
where $\mathcal{O}^{(h)}(g(m,n))$ is defined in Lemma 2.1.
It is worth noting that for closed manifolds with intrinsic dimension $d\leq 3$, the estimator in [landa2023robust] achieves a second-order error of $\mathcal{O}(h^{2})$—a result conjectured to hold for higher dimensions as well. In contrast, our estimator for manifolds with boundary is limited to first-order accuracy $\mathcal{O}(h)$. A primary cause of this limitation is that the mean curvature $H(\mathbf{x})$ induced by the boundary is unknown (see theorem 3.1), thereby preventing the implementation of higher-order corrections. Nevertheless, the proposed first-order estimator in Eq. (41) is computationally feasible, and is implemented using the estimated distances to the boundary from Algorithm 2 to compute the correction terms in $\hat{q}_{i}$.